comlex no *** 1+i2+i3+i4/1+i

comlex no
***  1+i2+i3+i4/1+i

Grade:12th pass

1 Answers

Nishant Vora IIT Patna
9 years ago
1+i2+i3+i4/1+i
i2=-1
i3=-i
i4=1

\frac{1+i^2+i^3+i^4}{1+i} = \frac{1+(-1)+(-i)+(1)}{1+i}=\frac{1-i}{1+i}
Now you can rationalize the denomenator